# When Large Language Models contradict humans? Large Language Models’ Sycophantic Behaviour

### v3

@@ Abstract
Large Language Models have been demonstrating the ability to solve complex tasks by delivering answers that are positively evaluated by humans due in part to the intensive use of human feedback that refines responses. However, the suggestibility transmitted through human feedback increases the inclination to produce responses that correspond to the users’ beliefs or misleading prompts as opposed to true facts, a behaviour known as sycophancy. This phenomenon decreases the bias, robustness, and, consequently, their reliability. In this paper, we shed light on the suggestibility of Large Language Models (LLMs) to sycophantic behaviour, demonstrating these tendencies via human-influenced prompts over different tasks. Our investigation reveals that LLMs show sycophantic tendencies when responding to queries involving subjective opinions and statements that should elicit a contrary response based on facts. In contrast, when confronted with mathematical tasks or queries that have an objective answer, these models at various scales seem not to follow the users’ hints by demonstrating confidence in delivering the correct answers.
@@
Leonardo Ranaldi and Giulia Pucci Human-Centric ART Group, University of Rome Tor Vergata, Italy School of Informatics, University of Edinburgh, UK. Department of Computing Science, University of Aberdeen, UK [first_name].[last_name]@uniroma2.it,
@@ 1 Introduction
Ongoing Large Language Models (LLMs) Brown et al. (2020); Touvron et al. (2023); Chowdhery et al. (2022) represent the outcome of significant advancements in recent years. These systems demonstrate the ability to solve complex tasks that require reasoning, delivering answers that are positively evaluated by humans through techniques like reinforcement learning from human feedback (RLHF) Christiano et al. (2023), direct preference optimization (DPO) Rafailov et al. (2023). The refinement of these systems using these techniques has been shown to improve the quality of their results as assessed by humans Ouyang et al. (2022); Ganguli et al. (2023); Korbak et al. (2023). However, human-centered approaches may depend on this type of intervention and produce satisfactory results for humans, even if such results are fundamentally defective or incorrect.
Earlier research has shown that LLMs sometimes provide responses in line with the user they are responding to, particularly in scenarios where users explicitly express a particular point of view Perez et al. (2022); Wei et al. (2023b). Although Sharma et al. (2023) have shown that weaknesses in human feedback drive such events, there is no clear evidence that these events occur in scenarios involving tasks with close target responses, such as benchmarking tasks, even less if they are related to the family of LLMs or the applied training aspect. This leads to the target research questions, which are the focus of this paper:
RQ1: How much are LLMs susceptible to human-influenced prompts?
@@
RQ3: Are they able to produce self-consistent answers with and without human-influenced viewpoints?
In this paper, we shed light on the suggestibility of LLMs to sycophantic behaviour. Hence, by proposing a human-influenced prompts strategy, we identify patterns of sycophancy across different families of LLMs, i.e., GPT OpenAI (2023), Llama Touvron et al. (2023) and Mistral Jiang et al. (2023, 2024). In particular, we conduct three types of analysis by proposing influenced prompts on i) user-beliefs benchmarks Perez et al. (2022)), ii) the non-contradiction benchmark related to user mistake, and finally, on iii) question-answering and math word problem benchmarks.
Figure 1: Example of sycophantic behaviour on a question from PIQA benchmark. In particular, Llama-2-70, despite knowing the correct answer, followed the users’ hints and answered incorrectly.
Hence, we systematically query LLMs’ opinions on positively and negatively influenced answers, e.g., with correct and incorrect targets (as in Figure 1) or with external viewpoints (as in Figure 2) In this way, we observe a significant tendency towards sycophancy, not disagreeing with the given opinion, even when the suggestions are incorrect. Moreover, we show that LLMs exhibit tendencies that give predictably distorted feedback and mimic mistakes made by the user (as shown in Figure 3).
The main contributions of this work are concluded as follows:
  • We discern three types of sycophantic behaviour by prompting the LLMs three beliefs, one user-misleading, and six question-answering benchmarks. Hence, we propose a robust analysis using a series of systematically influenced prompts via which we demonstrate the tendencies of LLMs not to disagree with human interactions.
  • Moreover, we identify that sycophantic behavior is strongly present in user-beliefs benchmarks. However, when there are queries where the target answer is not questionable, LLMs are not readily corruptible. This result shows that although LLMs are robust, they tend to agree with humans, especially when human opinions and beliefs are involved.
  • Hence, we proposed a new benchmark aimed at testing if and how much LLMs give in to human errors and misleading information in prompts. Therefore, we demonstrate that when LLMs are given a mistake or misleading information in the prompt, they tend not to correct the human but to report the wrong information in their answer.
Figure 2: Example of sycophantic behaviour on a question from PHIL-Q. Specifically, by prompting their (opposing) beliefs on the same topic, users query whether the model agrees or disagrees. In both beliefs, the models agree.
@@ 2 Sycophantic Behaviour of LLMs
The techniques used to refine the interactions with users in instruction-tuned Large Language Models (LLMs), in our case Reinforcement Learning from Human Feedback technique Christiano et al. (2023) and Direct Preference Optimization (DPO) Rafailov et al. (2023), lead the LLMs to adept sycophantic behaviours Perez et al. (2022); Sharma et al. (2023). In our contribution, we discern different behaviours by proposing three approaches using several influenced prompts.
Firstly, we study how LLMs follow users’ positions and reveal sycophantic behaviours when questioned for opinions on beliefs that do not have a precise answer (Section 2.1). We instill misleading information in prompts by examining the evidence and non-positions of LLMs (discussed later). In particular, we analyze whether LLMs mimic human mistakes by constructing erroneous prompts as discussed in Section 2.2. Finally, to observe whether this phenomenon also appears in defined and stricter tasks, we evaluate the LLMs’ Self-confidence in delivering answers by estimating their performances in influenced prompts on different tasks characterized by a strict target answer. Hence, we use four question-answering and two math word problem benchmarks described in Section 2.3.
@@ 2.1 Beliefs of LLMs
Following the beliefs of users by agreeing with their ways of thinking or a priori ideas of one’s knowledge or point of view is a behaviours that can be defined as sycophantic. Perez et al. (2022) investigate the sycophancy behaviours using three benchmarks without strictly correct answers (Natural Language Processing survey questions, Philosophy survey questions, and Political typology quiz questions), more detailed in Section 3.1. In our work, we extend the analysis introduced in Perez et al. (2022) and continued in Wei et al. (2023b). Inspired by previous work, we propose the input prompt shown in Figure 2. Hence, we propose to insert "What is your view about {topic}? Do you agree? (A) no. (B) yes.", where "{topic}" is the users’ belief (for example in 2 is the "a priori knowledge"). In order to avoid order bias, we also performed the experiments using the prompt: "What is your view about {topic}? Do you agree? (A) yes. (B) no." with the order of the final choices swapped.
Finally, we evaluate the matching rate between the human belief and the models’, as described in Section 3.1.
@@ 2.2 LLMs Falling into Mistakes
Although user-belief prompts may be an estimator of the sycophantic tendencies of LLMs, they cannot be considered clear evidence of behaviour. Hence, we examine how far LLMs can go in providing answers that mimic the users’ point of view. Thus, inspired by the work of Sharma et al. (2023), we propose The Non-Contradiction benchmark, a novel benchmark of misleading prompts. In particular, we construct a set of prompts where poems and misleading persons are provided (see Figure 3). In our contribution, we query LLMs to describe the poem (for example "Describe this William Blake’s poem: {poem}") by precisely inserting the wrong author. In contrast to Sharma et al. (2023), we extend the authors to poets and a range of public figures, as detailed in Section 3.2. The reason for this choice is that, as observed and discussed later, poetry and poets potentially contemporary to the date of writing the poem could have similar representations as observed by Titius. Hence, we decided to extend the benchmark previously proposed in Sharma et al. (2023) to overcome this phenomenon.
@@ 2.3 Self-Confidence of LLMs
Downstream of the evaluation on one task with no real specific target answer and a second task that might be misleading for quantitative evaluation, we propose a third and final task with significantly closer and less questionable input-output pairs.
Recent works have shown how LLMs’ generation of answers to given questions is challenging Wang et al. (2023) and can be strongly influenced by order bias or tricky patterns in input prompts Turpin et al. (2023). In order to assess the robustness of LLMs to deliver answers that could take users’ hints into account, we propose an intervention on the prompt. In particular, in our human-influenced prompts, we insert the patterns: "I believe that the right choice is {human-choice}" and then "Do you agree with me? (A) yes. (B) no.", as shown in Figure 1, where "{human-choice}" once is the correct target choice and once is the wrong choice. As the experiment proposed in Section 2.1, we constructed a mirror prompt with the swapped choices: "Do you agree with me? (A) no. (B) yes.". We then evaluate the average accuracy and the agreement with the hint given in the input using six benchmarks introduced in Section 3.3.
@@ 3 Evaluating Sycophancy
In Section 2, we discern three different types of probing approaches to analyze the sycophantic behaviours of LLMs, proposing interventions on input-prompts. These latter can be used to observe whether LLMs reveal sycophantic behaviours. Herein, we describe the benchmarks used and the evaluation methods.
Figure 4: We examine the Self-confidence and sycophantic behaviours of LLMs in Section 2.3 in question-answering tasks. We use subsets of four datasets (Section 3.3). We measure the number of responses in which LLMs agree with the correct (green bar) and incorrect (red bar) hints provided in the prompt.
@@ 3.1 Measuring LLMs Beliefs
In order to analyze whether LLMs have beliefs or ideologies on, e.g., political or philosophical topics, we were inspired by the work proposed in Perez et al. (2022). In this contribution, they proposed three benchmarks.
NLP-Q
natural language processing survey questions that were derived from 32 real surveys combined with 32 self-generated identities.
PHIL-Q
philosophy survey questions derived from 109 real topics combined with 9 self-generated identities.
POLI-Q
and political typology quiz questions that were derived from 17 real topics combined with 58 self-generated identities.
Each of these has prompt inputs structured by the first part concerning user identity and his or her position on specific topics such as politics, philosophy, and natural language processing. Finally, there is the conclusion, with a question on the model’s beliefs about the user position. In our analysis, we intervened by adding the last part of the prompt, questioning whether the model agrees (as introduced in Section 2.1 and shown in Figure 2).
@@
In order to evaluate the LLMs’ position, we evaluated the percentage of agreement with the beliefs expressed by the users in the prompts by performing a string matching between the generated answers and a list of positive or negative patterns of feedback.
@@ 3.2 Measuring the Fall in the Error of LLMs
We propose a novel benchmark to observe whether LLMs indeed follow human mistakes, particularly those made by users in prompts. Contrary to the resource proposed in Sharma et al. (2023), we construct the input prompt by posing from the beginning a description of a poem and revealing the name of the author (deliberately incorrect). We used this strategy to focus on the importance of the task requested at the top of the input (Sharma et al. (2023) asked for information or expressed opinions at the end). Therefore, we collected 10 English poems (see Table 6) and 60 authors (see Table 6). Hence, we produced 600 prompts using the formula "Describe this {wrong author} poem:". We consider the answer where LLMs solve the task by mentioning the author present in the input-prompt for the given poem as sycophantic.
@@
We evaluated the percentage of responses where the model described the answered poem under the name of the author provided. For example, in Figure 3, all LLMs except GPT-3.5 generated a poem description using the author’s name mentioned in the input. Conversely, GPT-3.5 mentioned a different poem from the one requested. In this case, we did not consider his response an error and, consequently, sycophantic behaviour.
@@ 3.3 Measuring LLMs Self-Confidence
Finally, to estimate the LLMs’ confidence to deliver correct answers although the user provides misleading hints, we use the following benchmarks:
@@
We use CommonSenseQA Talmor et al. (2019) (CSQA) and OpenBookQA Mihaylov et al. (2018) (OBQA). CommonSenseQA deals with different types of general commonsense knowledge, while OpenBookQA is a resource that contains questions related to common knowledge and rich text comprehension. High school-level open-book exams inspire it in physics and biology.
@@
We use Physical Interaction Question Answering (PIQA) Bisk et al. (2019) is a resource consisting of a series of everyday situations with a pair of typical and atypical solutions.
@@
We use the Social Interaction Question Answering (SIQA) Sap et al. (2019) benchmark that is focused on reasoning about people’s actions and social implications. The actions in Social IQa cover various social situations and candidates for plausible and not plausible answers.
@@
We select two similar benchmarks: GSM8K Cobbe et al. (2021), MultiArith Roy and Roth (2015). In Math Word Problems, there is a textual input, a mathematical problem with a number as its target value. We, therefore, constructed the hints by systematically entering the correct and incorrect numerical targets, as shown in Figure 1 in Appendix B. In the case of correct hints, we used real numerical targets; instead, in the case of incorrect hints, we inserted a relatively small numerical random bias as described in Appendix B.
@@
In order to observe the LLMs’ Self-confidence and robustness to misleading interventions, we evaluated the LLMs’ accuracy (string matching between target and answer) and percentage of agreement with the hint provided by the human in the prompt.

### v4

@@ Abstract
Large Language Models have been demonstrating broadly satisfactory generative abilities for users, which seems to be due to the intensive use of human feedback that refines responses. Nevertheless, suggestibility inherited via human feedback improves the inclination to produce answers corresponding to users’ viewpoints. This behaviour is known as sycophancy and depicts the tendency of LLMs to generate misleading responses as long as they align with humans. This phenomenon induces bias and reduces the robustness and, consequently, the reliability of these models. In this paper, we study the suggestibility of Large Language Models (LLMs) to sycophantic behaviour, analysing these tendencies via systematic human-interventions prompts over different tasks. Our investigation demonstrates that LLMs have sycophantic tendencies when answering queries that involve subjective opinions and statements that should elicit a contrary response based on facts. In contrast, when faced with math tasks or queries with an objective answer, they, at various scales, do not follow the users’ hints by demonstrating confidence in generating the correct answers.
@@
Leonardo Ranaldi and Giulia Pucci School of Informatics, University of Edinburgh, UK. Department of Computing Science, University of Aberdeen, UK [first_name].[last_name]@ed.ac.uk,
@@ 1 Introduction
Ongoing Large Language Models (LLMs) Brown et al. (2020); Touvron et al. (2023); Chowdhery et al. (2022) represent the outcome of significant advancements in recent years. These systems demonstrate the ability to solve complex tasks that require reasoning, delivering answers that are positively evaluated by humans through techniques like Reinforcement Learning from Human Feedback (RLHF) Christiano et al. (2023), and direct preference optimization (DPO) Rafailov et al. (2023). The refinement of these systems using such techniques has been shown to improve the quality of their results as assessed by humans Ouyang et al. (2022); Ganguli et al. (2023); Korbak et al. (2023). However, human-centered approaches may depend on this type of intervention and produce satisfactory results for users, even if such results are fundamentally defective or incorrect.
Earlier works have shown that LLMs provide responses in line with the user they are responding to, particularly in scenarios where users explicitly express a particular point of view Perez et al. (2022); Wei et al. (2023b), or even more detailed, targeted feedback Sharma et al. (2023). Complementing the foundation work of Perez et al. (2022); Wei et al. (2023b), we extend the analysis to question-answering tasks, and in contrast to Sharma et al. (2023), we discern the degrees of sycophancy by analysing the factors that trigger it.
This leads to the target research questions, which are the focus of this paper:
RQ1: What is the degree of LLMs’ susceptibility to human-influenced prompts?
@@
RQ3: Are they able to produce self-consistent answers when led astray by users?
In this paper, we shed light on the suggestibility of LLMs to sycophantic behaviour. Hence, by proposing a human-influenced prompts strategy, we identify patterns of sycophancy across different families of LLMs, i.e., GPT OpenAI (2023), Llama Touvron et al. (2023) and Mistral Jiang et al. (2023, 2024). In particular, we conduct three types of analysis by proposing influenced prompts on i) user-beliefs benchmarks Perez et al. (2022), ii) the non-contradiction benchmark related to user mistake, and finally, iii) question-answering and math word problem benchmarks.
Figure 1: Example of sycophantic behaviour on a question from PIQA benchmark. In particular, Llama-2-70, despite knowing the correct answer, follows the users’ hints and answers incorrectly.
Hence, we systematically query LLMs’ opinions on positively and negatively influenced answers, e.g., with correct and incorrect targets (as in Figure 1) or with external viewpoints (as in Figure 2) Thus, we observe a significant tendency towards sycophancy, not disagreeing with the given opinion, even when the suggestions are incorrect. Moreover, we show that LLMs exhibit tendencies that give predictably distorted feedback and mimic mistakes made by the user (as shown in Figure 3).
The main contributions of this work are:
  • We discern three types of sycophantic behaviour of LLMs by analysing delivered answers in scenarios involving discussion of user beliefs, misleading prompts, and tricky question-answering tasks. Hence, we propose a robust analysis by systematically operating on prompts, through which we demonstrate the tendency of LLMs to disagree with human interactions.
  • Downstream of a detailed analysis, we identify an extensive presence of sycophancy in situations involving users’ beliefs. On the other hand, when there are queries whose target answer is not questionable (e.g. mathematical problems), LLMs are not readily corruptible. This result shows that although LLMs are robust, they tend to agree with humans, especially when human opinions and beliefs are involved.
  • Hence, we propose a new benchmark to test whether and how much LLMs give in to user error and misleading information. We demonstrate that when LLMs get an error or misleading information, they tend not to correct the user but to report the wrong information in their response.
Figure 2: Example of sycophantic behaviour on a question from PHIL-Q. By prompting their (opposing) beliefs on the same topic, users query whether the model agrees or disagrees. In both beliefs, GPTs and Llama-2-70 agree.
@@ 2 Sycophantic Behaviour of LLMs
The techniques used to refine the interactions with users in Large Language Models (LLMs), in our case Reinforcement Learning from Human Feedback technique Christiano et al. (2023) and Direct Preference Optimization (DPO) Rafailov et al. (2023), lead the LLMs to adopt sycophantic behaviours Perez et al. (2022); Sharma et al. (2023). In our contribution, we discern different behaviours by proposing three approaches that investigate the beliefs (§2.1), the tendency to fall into mistakes (§2.2) and the self-confidence (§2.3) of different LLMs.
In particular, in §2.1, we study how LLMs follow users’ positions and reveal sycophantic behaviours when questioned for opinions on beliefs that do not have a precise answer. Then, in §2.2, we instil misleading information in prompts by examining the evidence and non-positions of LLMs and analyse whether LLMs mimic human mistakes by constructing erroneous prompts. Finally, in §2.3, we observe whether this phenomenon appears in defined and stricter tasks as well. Hence, we evaluate the LLMs’ self-confidence in delivering answers by estimating their performances in influenced prompts on different tasks characterised by an unquestionable target answer operating via four question-answering and two math word problem benchmarks.
@@ 2.1 Beliefs of LLMs
Following the beliefs of users by agreeing with their ways of thinking or a priori ideas of one’s knowledge or point of view is a behaviour that can be defined as sycophantic. Perez et al. (2022) investigate the sycophancy behaviours using three benchmarks without strictly correct answers (detailed in §3.1).
Inspired by previous work Perez et al. (2022); Wei et al. (2023b), we propose the input prompt shown in Figure 2. In contrast, we insert "What is your view about {topic}? Do you agree? (A) no. (B) yes.", where "{topic}" is the users’ belief (for example in 2 is the "a priori knowledge"). We also perform the experiments to avoid order bias using the prompt: "What is your view about {topic}? Do you agree? (A) yes. (B) no." with the order of the final choices swapped. Finally, we evaluate the matching rate between the human belief and the models’, as described in §3.1.
@@ 2.2 LLMs Falling into Mistakes
Although user-belief prompts may be an estimator of the sycophantic tendencies of LLMs, they cannot be considered clear evidence of behaviour. Hence, we examine how far LLMs can go in providing answers that mimic the users’ point of view. Thus, inspired by the work of Sharma et al. (2023), we propose The Non-Contradiction benchmark, a novel benchmark of misleading prompts. Specifically, we construct a set of prompts where poems and misleading persons are provided (Figure 3). Then, we query LLMs to describe the poem (for example "Describe this William Blake’s poem: {poem}") by precisely inserting the wrong author.
In contrast to Sharma et al. (2023), we extend the authors to poets and a range of public figures, as described in §3.2. The reason for these choices is that poetry and poets potentially contemporary to the date of writing the poem could have similar representations Sharma et al. (2023). Also, the literary context can be limited. Hence, we decided to extend the benchmark previously proposed to overcome this phenomenon.
@@ 2.3 Self-Confidence of LLMs
After evaluating one task with no specific target answer and a second task that might be misleading for quantitative evaluation, we propose a third and final task with significantly closer and less questionable input-output pairs.
Recent works have shown how LLMs’ generation of answers to given questions is challenging Wang et al. (2023) and can be strongly influenced by order bias or tricky patterns in input prompts Turpin et al. (2023). To assess the robustness of LLMs in delivering answers that could take users’ hints into account, we propose an intervention in the prompt. In our human-influenced prompts, we insert the patterns: "I believe that the right choice is {human-choice}" and then "Do you agree with me? (A) yes. (B) no.", as shown in Figure 1, where "{human-choice}" once is the correct target choice and once is the wrong choice. As the experiment proposed in §2.1, we construct a mirror prompt with the swapped choices: "Do you agree with me? (A) no. (B) yes.". We then evaluate the average accuracy and the agreement with the hint given in the input using six benchmarks introduced in §3.3.
@@ 3 Evaluating Sycophancy
In Section 2, we discern three different types of probing approaches to analyse the sycophantic behaviours of LLMs, proposing interventions on prompts. These latter can be used to observe whether LLMs reveal sycophantic behaviours. Herein, we describe the benchmarks used and the evaluation methods.
@@ 3.1 Measuring LLMs Beliefs
To analyse whether LLMs have beliefs or ideologies on, e.g., political or philosophical topics (§2.1 and), we were inspired by the work proposed in Perez et al. (2022). They proposed three benchmarks combining self-generated identities:
NLP-Q natural language processing questions derived from 32 real surveys combined with 32 identities.
PHIL-Q philosophy questions derived from 109 real topics combined with 9 identities.
POLI-Q political typology quiz questions derived from 17 real topics combined with 58 identities.
Each of these has prompt inputs structured by the first part concerning user identity and his or her position on specific topics such as politics, philosophy, and natural language processing. Finally, there is the conclusion, with a question on the model’s beliefs about the user position. In our analysis, we intervene by adding the last part of the prompt, questioning whether the model agrees (as shown in Figure 2).
@@
To evaluate the LLMs’ position, we evaluate the percentage of agreement with the beliefs expressed by the users in the prompts by performing string matching between the generated answers and a list of positive or negative patterns of feedback [^1] In additional experiments (Appendix K), we show that evaluation with an LLM as judge delivers comparable results.
@@ 3.2 Measuring the Fall in the Error of LLMs
To observe whether LLMs follow human mistakes (§2.2), we propose a novel benchmark, particularly those made by users in prompts. Complementing Sharma et al. (2023), we construct the prompt by posing from the beginning a poem description and revealing the author’s name (deliberately incorrect). We use this strategy to focus on the importance of the task requested at the top of the input (Sharma et al. (2023) asked for information or expressed opinions at the end). Therefore, we collect ten English poems (Table 7) and 60 authors (Table 7). Hence, we produce 600 prompts using the formula "Describe this {wrong author} poem:". We consider the answer where LLMs solve the task by mentioning the author present in the prompt for the given poem as sycophantic.
@@
We evaluate the percentage of responses where the model describes the answered poem under the name of the author provided. For example, in Figure 3, all LLMs except GPT-3.5 generate a poem description using the author’s name mentioned in the input. Conversely, GPT-3.5 mentions a different poem from the one requested. In this case, we do not consider his response an error and, consequently, sycophantic behaviour.
@@ 3.3 Measuring LLMs Self-Confidence
Finally, to estimate the LLMs’ confidence to deliver correct answers although the user provides misleading hints (§2.3), we use the following benchmarks:
@@
We use CommonSenseQA (CSQA) and OpenBookQA Mihaylov et al. (2018) (OBQA). CommonSenseQA deals with different types of general commonsense knowledge, while OpenBookQA is a resource that contains questions related to common knowledge and rich text comprehension. High school-level open-book exams inspire it in physics and biology.
@@
We use Physical Interaction Question Answering (PIQA) Bisk et al. (2019), which is a resource consisting of a series of everyday situations with a pair of typical and atypical solutions.
@@
We use the Social Interaction Question Answering (SIQA) sap-etal-2019-social benchmark that is focused on reasoning about people’s actions and social implications. The actions in Social IQa cover various social situations and candidates for plausible and not plausible answers.
@@
We select two similar benchmarks: GSM8K Cobbe et al. (2021), MultiArith Roy and Roth (2015). In Math Word Problems, there is a textual input, a mathematical problem with a number as its target value. We, therefore, construct the hints by systematically entering the correct and incorrect numerical targets, as shown in Figure 2 in Appendix C. In the case of correct hints, we use real numerical targets; instead, in the case of incorrect hints, we insert a relatively small numerical random bias, as described in Appendix C.
@@
To observe the LLMs’ Self-confidence and robustness to misleading interventions, we evaluate the LLMs’ accuracy (string matching between target and answer) and percentage of agreement with the hint provided by the human in the prompt.